OpenStudy (anonymous):

Please Help Use the properties of logarithms to write the following expression as a single logarithm: 2logx+log5y-1/2logz

OpenStudy (solomonzelman):

not sure....

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

I do have an example though. http://answers.yahoo.com/question/index?qid=20100827201107AAVGKmI

OpenStudy (tkhunny):

What have you tried? 2logx+log5y-(1/2)logz Take care of those external coefficients. \(\log(x^{2}) + \log(5y) - \log(\sqrt{z})\) Now what?

OpenStudy (anonymous):

right now I just have,